you obtained meet typical vitamh anu e eerts 2. A liter of a TPN solution contains 500 millWiters of 50 percer dextrose solution and 500 mililiters of 5 percent anino acd solution. Determine the daily energy and protein ntakes of CH APİER SIXTE
Answer #1

1. 500ml in 50% of dextrose:

Usually 10% of dextrose contains= 10g of glucose/100ml

So 50% dextrose contains = 50g of glucose / 100ml

Then for 500ml= 5× 50g of glucose= 250g/ 500ml

Patient is getting 2 lit= 2000ml =250×4= 1000g of glucose

1g of glucose gives= 4kcal,  

1000g of glucose = 1000× 4= 4000kcal.

Patient is getting 4000kcal energy and no protein from dextrose.

2. 500ml TPN contains 5% amino acids,

In 100ml / 5% aminoacid gives 5 gram protein

So in 500ml = 5× 5 grams= 25 grams protein.

Patient is getting 2 liter,

So in 2000ml = 25× 4 grams= 100g of protein

So patient is getting 100 gram of protein per day.